OpenStudy (anonymous):

Medal The oil sands reserves of Canada are estimated at 3.5 × 1011 tonnes (t). How long (in years) would this supply last at a rate of consumption of 8.4 × 108 tonnes per year? Retain the proper number of digits in your answer. Time = years

OpenStudy (anonymous):

\[\huge \frac{3.5\times 10^{11}}{8.4\times 10^8}\]

OpenStudy (sidmanfu):

do you still need help?